OpenStudy (anonymous):

Essay; show all work. A simple dartboard has three areas… the main board has a radius of 11 inches, there is a circle with a radius of 5 inches, and the bullseye has a radius of 2 inches. What is the probability of a random dart landing inside the bulls eye? Round to the nearest thousandth. (Hint: Use A= Pie r2

OpenStudy (anonymous):

it is the area of the bulls eye divided by the area of the dart board

OpenStudy (anonymous):

\[\frac{2^2\pi}{11^2\pi}=\frac{4}{121}\]